# Accelerating Large Language Model Pretraining via LFR Pedagogy: Learn, Focus, and Review

### v1

@@ Abstract
Large Language Model (LLM) pretraining traditionally relies on autoregressive language modeling on randomly sampled data blocks from web-scale datasets. We take inspiration from human learning techniques like spaced repetition to hypothesize that random data sampling for LLMs leads to high training cost and low quality models which tend to forget data. In order to effectively commit web-scale information to long-term memory, we propose the LFR (Learn, Focus, and Review) pedagogy, a new dynamic training paradigm which focuses and repeatedly reviews complex data blocks at systematic intervals based on the model’s learning pace and progress. LFR records the model perplexities for different data blocks and frequently revisits blocks with higher perplexity which are more likely to be forgotten. We pretrain the GPT-2 models (124M - 1.5B) from scratch on the OpenWebText dataset using LFR. We test on downstream tasks from the language modeling, question answering, translation, and problem solving domains to achieve consistently lower perplexity and higher accuracy than the baseline OpenAI models, while obtaining a 20x pretraining speed-up.
@@ 1 Introduction
LLMs have achieved remarkable success in understanding and generating human language. This success is driven by the ever-increasing model parameter sizes which require web-scale training datasets like OpenWebText Radford et al. (2019); ope , CommonCrawl Penedo et al. (2023); Raffel et al. (2023), and Pile Gao et al. (2020), leading to unsustainable training costs. For example, pretraining the GPT-4 model OpenAI et al. (2024) was estimated to have cost around \$100M USD over a period of 3-4 months using 25k NVIDIA A100 GPUs gpt . As such, a key challenge for unlocking the next generation of language models is to significantly reduce training costs while retaining or improving downstream task performance.
Popular LLMs, like the GPT series Radford et al. (2019); Brown et al. (2020); OpenAI et al. (2024), Llama series Touvron et al. (2023, 2023); lla (2024), Mistral Jiang et al. (2023), and others, were trained utilizing autoregressive/causal techniques, where the goal of training is to generate a plausible next token based on a randomly sampled input data block from the training corpus. However, data scraped from the web is highly redundant and presents varying complexity. For example, Table 1 shows two randomly selected example texts from the OpenWebText Radford et al. (2019); ope dataset. If a human is asked to learn these two texts, they might spend more time learning the details in the first block of text than those in the second. The first block contains more complex information with facts and dates, while the second block contains anecdotal information, which can be understood faster. They might read the first example more times than the second. However, autoregressive language modeling does not mimic such behavior.
The slave rebellion of the early 70s BCE led by Spartacus, the Thracian gladiator, would present Crassus, made praetor in 73 BCE, with an opportunity to flex his military muscle and gain further prestige with the Roman people. The slave army, numbering between 70,000 and 120,000, was a serious threat and had already defeated two separate Roman armies and two consuls. Now they were ravaging the southern Italian countryside and it was Crassus who was entrusted with finally removing this thorn from Rome’s heel.
Sata, who is older than Japan’s nearly 60-year-old civilian nuclear industry, remembers a time without nuclear power. Families were allowed just a few lightbulbs in the 1940s, because the electrical system was still in its infancy. “There was a TEPCO office in the neighborhood,” she recalls, and when a bulb burned out, “we had to bring it there” to trade it for a new one. The advent of nuclear power meant that the Japanese could consume as much electricity as they wanted.
Table 1: Examples of Text Blocks in the OpenWebText Dataset
@@
    Profile LLM pretraining to observe multiple descent behaviour for 25% of the training tokens that are forgotten multiple times.
@@
    Pretrain GPT-2 models (124M-1.5B) on the OpenWebText dataset with 4 AMD MI210 (345M-1.5B) and 4 AMD MI100 GPUs (124M), where the 1.5B model training took less than 20 days, and evaluate them on six downstream tasks covering language modeling, question answering, translation, and problem solving. In fact, we believe that we are the first to train the GPT-2 models on a cluster of AMD GPUs, demonstrating an alternative GPU platform for LLM training.
@@
    LFR results in significantly lower perplexity and higher accuracy compared to baseline OpenAI models, achieving these improvements with 20x fewer training iterations.
@@
    Observe that LLMs first learn conversational and anecdotal data, before being able to retain factual and instructional information in long-term memory.
@@ 2 Related Work
Prior works on efficient pretraining of LLMs using data selection have primarily focused on using distance metrics and clustering techniques. Tirumala et al. (2023) proposes D4, which deduplicates and selects cluster centers in the embedding space generated by pretrained models. SemDeDup Abbas et al. (2023) prunes semantic duplicates using pretrained models. It can successfully prune 50% of the training data with minimal performance loss. MiniPile Kaddour (2023); min uses the pretrained E5-Large Wang et al. (2024) model to embed documents in the Pile dataset and clusters them to select a smaller pretraining corpus of $\sim$6GB. DSIR Xie et al. (2023) proposes selecting subsets from large unlabeled datasets through importance resampling to match the distribution of the target dataset. However, considering the high cost of training, it is unsustainable to sample a new subset and pretrain the LLM from scratch for every new downstream task. While these works address the common redundancy issue present in LLM pretraining corpuses, they face three major drawbacks. First, they require pretrained models for calculating the distance metric on embeddings. This step must ensure that the embeddings found by the pretrained model are similar to those of the target model to ensure the transferability of sample importance. Second, clustering-based methods do not scale to encompass the diversity of data in web-scale datasets. This can lead to a loss in downstream task accuracy (see Section 5). Third, they are slow and can take as long as 2-3 hours to generate 6GB subsets of the OpenWebText dataset (38GB) as shown in Section 5.
@@ 3.1 LLM Pretraining Objective
Given an LLM model parameterized by weights $\theta$ and a web-scale dataset $D$, we first tokenize all the documents in the dataset and obtain context-length sized sequences of tokens $s_{i}$ such that the training corpus becomes $D=\{s_{1},s_{2},s_{3},...s_{n}\}$. For the OpenWebText dataset used in this paper, the context-length is 1024 and there are a total of 9B training tokens. Given one such sequence of tokens $s_{i}=\{x_{1},x_{2},...x_{n}\}$, the training objective is to autoregressively predict the next $M$ tokens:
@@ 3.2 Observations from Training on Randomly Sampled Data
In order to better understand the drawbacks of this traditional training technique, we profile the pretraining process for the 345M parameter GPT-2 model. The training hyperparameters are provided in Appendix A.1. Similarly to Marion et al. (2023), we use perplexity as a metric to monitor the training progress. Given a token sequence $s_{i}=\{x_{1},x_{2},...,x_{n}\}$ from the dataset $D$, perplexity is computed as:
@@
To generate profiling information, we record the data block ID (each data block consists of 1024 tokens) and model perplexity each time a block is randomly sampled from the training corpus across all training iterations. At the end of the training process, we have a record of the perplexities over time for each data block. Since we train the GPT-2 345M model for 8 epochs, each data block has at least 8 perplexity values captured at different training iterations. From this record, we classify each data block as one of the following:
@@
We observe that 25% of data blocks are forgotten at least once during training and that of the data blocks that are forgotten, 82% are forgotten multiple times during training, i.e., they display multiple descent behavior (Figure 1). Xia et al. (2022) reported a double-descent behavior for the OPT models Zhang et al. (2022), and our above experiment further demonstrates that the “forgotten” can happen multiple times in LLM pretraining.
Figure 1: Perplexities of data samples being forgotton by the GPT2-345M model.
@@ 4 LFR Training Methodology
Based on our profiling observations in Section 3.2 we propose to replace traditional autoregressive language modeling methods with Spaced Repetition Tabibian et al. (2019). Spaced Repetition is an evidence-based learning method proven to improve information retention and learning pace in humans Smolen et al. (2016). In this technique, challenging pieces of information are reviewed more often, at regular intervals, and easier pieces of information are presented to the learner less often. We pretrain our models with a combination of the following three steps:
@@
    Learn: Initially, we present the model with the entire dataset and train on randomly selected data blocks for $p_{1}$ epochs, as normally seen in the traditional approach. $p_{1}$ can be configured by the user based on the available compute budget, model, and dataset. During the $p_{1}$ epochs, we also record the perplexities for all data blocks. The next 2 phases can be repeated $reps$ number of times based on the available compute budget.
@@
    Focus: From the collected perplexities from the previous phase, we discard $s_{1}$of the data blocks corresponding with the lowest perplexities. In doing so, we provide a mechanism for shifting the model’s focus towards learning data blocks that were determined to be difficult. In this phase of training, we restrict the random sampling of data points to this reduced subset for $p_{2}$ epochs. $s_{1}$ and $p_{2}$ are user-controlled hyperparameters.
@@
    Review: Next, we reintroduce all of the removed data blocks and train the model on the entire corpus for $p_{3}$ epochs (reconfigurable). This ensures that we allow the model to review and revisit data blocks which it may have forgotten. We record the perplexities for all data blocks again in this training phase.
Our algorithm is detailed in Algorithm 1. For pretraining the GPT-2 models, we choose the following combination of the aforementioned steps:

### v2

@@ Abstract
Traditional Large Language Model (LLM) pretraining relies on autoregressive language modeling with randomly sampled data from web-scale datasets. Inspired by human learning techniques like spaced repetition, we hypothesize that random sampling leads to high training costs, lower-quality models, and significant data forgetting. To address these inefficiencies, we propose the Learn-Focus-Review (LFR) paradigm—a dynamic training approach that adapts to the model’s learning progress. LFR tracks the model’s learning performance across data blocks (sequences of tokens) and prioritizes revisiting challenging regions of the dataset that are more prone to being forgotten, enabling better retention and more efficient learning. Using the LFR paradigm, we pretrained Llama and GPT models on the SlimPajama and OpenWebText datasets, respectively. These models were evaluated on downstream tasks across various domains, including question answering, problem-solving, commonsense reasoning, language modeling, and translation. Compared to baseline models trained on the full datasets, LFR consistently achieved lower perplexity and higher accuracy, while using only 5%–19% of the training tokens. Furthermore, LFR matched the performance of industry-standard Pythia models with up to 2$\times$ the parameter count, using just 3.2% of the training tokens, demonstrating its effectiveness and efficiency.
@@ 1 Introduction
LLMs have achieved remarkable success in understanding and generating human language. This success is driven by the ever-increasing model parameter sizes which require web-scale training datasets like SlimPajama Soboleva et al. (2023), OpenWebText Radford et al. (2019); ope , CommonCrawl Penedo et al. (2023); Raffel et al. (2023), and Pile Gao et al. (2020), leading to unsustainable training costs. For example, pretraining the GPT-4 model OpenAI et al. (2024) was estimated to have cost around \$100M USD over a period of 3-4 months using 25k NVIDIA A100 GPUs gpt . As such, a key challenge for unlocking the next generation of language models is to significantly reduce training costs while retaining or improving downstream task performance.
Popular LLMs, like the GPT series Radford et al. (2019); Brown et al. (2020); OpenAI et al. (2024), Llama series Touvron et al. (2023); lla (2024), Gemini Team (2024), Mistral Jiang et al. (2023), and others, were trained utilizing autoregressive/causal techniques, where the goal of training is to generate a plausible next token based on a randomly sampled input data block from the training corpus. However, data scraped from the web is highly redundant and presents varying complexity. For example, Table 1 shows two randomly selected example texts from the SlimPajama Soboleva et al. (2023) dataset. The SlimPajama dataset contains text from a variety of sources such as CommonCrawl, Github, ArXiv, Wikipedia, etc. If a human is asked to learn these two texts, they might spend more time learning the details and semantics of the code in the second block of text than the text in the first block. The second block contains more complex information with previously unseen code semantics, while the first block contains anecdotal information about a bakery, which can be understood faster. They might read the second example more times than the first. However, autoregressive language modeling does not mimic such behavior.
"Bread made to order and delivered fresh to you from my micro-bakery. All ingredients are seasonal and locally sourced." Find Rise ’n’ Shine Bakehouse regularly at our Emsworth Farmers’ Market. Find Foccacia, bagels, baguettes, french toast, sourdough and much more! Which markets are we attending?
NAMESPACE_PH0LY_BEGIN(thread) class PH0LY CriticalSection public: CriticalSection(); CriticalSection(); CriticalSection(const CriticalSection&\&&); CriticalSection&\&& operator = (const CriticalSection&\&& rhs); void Enter(); void Leave(); private: CRITICAL_SECTION m_cs; ; NAMESPACE_PH0LY_END
Table 1: Examples of Text Blocks in the SlimPajama Dataset. The second block is considerably hard to learn.
Figure 1: LLM training timeline
@@
    Profile LLM pretraining to observe multiple descent behaviour in 25-78% of the training tokens, which are forgotten multiple times during training.
@@
    Pretrain the Llama-2 models (120M-500M) using 4 NVIDIA A6000s and and the GPT-2 models (124M-1.5B) using 4 AMD MI210 and 4 MI100 GPUs from scratch on the SlimPajama and OpenWebText datasets and evaluate them on various downstream tasks.
@@
    LFR results in significantly lower perplexity and higher accuracy compared to baseline models trained on the full dataset using random sampling, achieving these improvements with 5.1$\times$ and 20$\times$ fewer training iterations for the Llama and GPT-2 models respectively.
@@
    Observe that LLMs first learn conversational and anecdotal data, before being able to retain factual, instructional, and coding language information in long-term memory. Furthermore, we compare our models with Pythia models of different scales trained on 30$\times$ more training tokens, and achieve comparable performance.
@@ 2 Related Work
Prior works on efficient pretraining of LLMs using data selection have primarily focused on using distance metrics and clustering techniques. Tirumala et al. (2023) proposes D4, which deduplicates and selects cluster centers in the embedding space generated by pretrained models. SemDeDup Abbas et al. (2023) prunes semantic duplicates using pretrained models. It can successfully prune 50% of the training data with minimal performance loss. MiniPile Kaddour (2023); min uses the pretrained E5-Large Wang et al. (2024) model to embed documents in the Pile dataset and clusters them to select a smaller pretraining corpus of $\sim$6GB. DSIR Xie et al. (2023) proposes selecting subsets from large unlabeled datasets through importance resampling to match the distribution of the target dataset. However, considering the high cost of training, it is unsustainable to sample a new subset and pretrain the LLM from scratch for every new downstream task. Lin et al. (2024) proposes a high-performance selection methodology, but suffers from dependence on a reference model for selection.
While these works address the common redundancy issue present in LLM pretraining corpuses, they face three major drawbacks. First, they require pretrained models for calculating the distance metric on embeddings. This step must ensure that the embeddings found by the pretrained model are similar to those of the target model to ensure the transferability of sample importance. Second, static clustering-based methods, which cluster data only at the start of training, fail to adapt to the diverse and dynamic nature of web-scale datasets. These methods overlook the fact that the difficulty of specific data points can vary across different training epochs, limiting their ability to effectively address the evolving learning needs of the model. This can lead to a loss in downstream task accuracy (see Section 5).
@@
Figure 2: PPL trajectories of data samples from the SlimPajama dataset as processed by the Llama-300M model, focusing on a subset of 50 samples for clarity. Notably, 78.5% of the samples exhibit this behavior, characterized by multiple descent patterns rather than a steady decline. This indicates that the model frequently forgets and relearns data during training, highlighting inefficiencies in traditional training dynamics
Figure 3: PPLs of data samples being forgotton by the GPT2-345M model. This multi-descent behavior is exhibited by 20% of the data.
@@ 3.1 LLM Pretraining Objective
Given an LLM model parameterized by weights $\theta$ and a web-scale dataset $D$, we first tokenize all the documents in the dataset and obtain context-length sized sequences of tokens, called data blocks, $s_{i}$ such that the training corpus becomes $D=\{s_{1},s_{2},s_{3},...s_{n}\}$. For the SlimPajama and OpenWebText datasets used in this paper, the context length is 1024 tokens, with a total of 627B and 9B tokens, respectively. Given one such sequence of tokens or data block, $s_{i}=\{x_{1},x_{2},...x_{n}\}$, the training objective is to autoregressively predict the next $M$ tokens:
@@ 3.2 Observations from Training on Randomly Sampled Data
In order to better understand the drawbacks of this traditional training technique, we profile the pretraining process for the 300M parameter Llama-2 model (single epoch training) and the 345M parameter GPT-2 model (multi-epoch training). The training hyperparameters and model configurations are provided in Appendix A.1. Similarly to Marion et al. (2023), we use perplexity as a metric to monitor the training progress. Given a token sequence $s_{i}=\{x_{1},x_{2},...,x_{n}\}$ from the dataset $D$, perplexity is computed as:
@@
The Llama models are trained with the SlimPajama dataset Soboleva et al. (2023) for 100k steps where we measure and plot the $PPL$ associated with the training set. In the case of the GPT-2 model, we train it for 8 epochs such that each data point in the OpenWebText dataset has at least 8 PPL values captured at different training iterations. The observed PPL values associated with each data block are classified as one of the following:
@@
Based on this classification, we observe that 78.5% of the data blocks are forgotten at least once in the Llama-2 model (Figure 2), compared to 25% in the GPT-2 model (Figure 3). We hypothesize that more data blocks are frequently forgotten in the Llama model due to the higher complexity and challenge posed by the SlimPajama dataset, as opposed to the OpenWebText dataset. It is important to note that the SlimPajama dataset is an aggregation of seven datasets, including sources such as GitHub, Wikipedia, and CommonCrawl. In fact, of the data blocks that are forgotten, 82% are forgotten multiple times during training, i.e., they display multiple descent behavior (Figure 3). Xia et al. (2022) reported a double-descent behavior for the OPT models Zhang et al. (2022), and our above experiment further demonstrates that the “forgetting” can happen multiple times in LLM pretraining.
@@ 4 LFR Training Methodology
Based on our profiling observations in Section 3.2 we propose to replace traditional autoregressive language modeling methods with Spaced Repetition Tabibian et al. (2019). Spaced Repetition is an evidence-based learning method proven to improve information retention and learning pace in humans Smolen et al. (2016). In this technique, challenging pieces of information are reviewed more often, at regular intervals, and easier pieces of information are presented to the learner less often. Our algorithm is detailed in Algorithm 1. We pretrain our models with a combination of the following three steps:
@@
    Learn: Initially, we present the model with the entire dataset and train on randomly selected data blocks for $p_{1}$ steps, as normally seen in the traditional approach (line 1 in Alg. 1). $p_{1}$ can be configured by the user based on the available compute budget, model, and dataset. In single-epoch training (lines 3-7 in Alg. 1), we measure the perplexities (PPLs) of all data samples in the training set and cluster the data embeddings (inputs to the model’s last layer). For multi-epoch training (lines 8-11 in Alg 1), we record the perplexities for all data blocks during the $p_{1}$ steps. Depending on the training style (single or multi-epoch), we either pass the clustered embeddings and PPL values or the PPL values observed during training to the next step. The following two phases can be repeated up to $reps$ times, depending on the available compute budget.
@@
    Focus: We provide two variations of the Focus stage based on the number of training epochs.
    1. (a)
      Single-epoch training: We discard $s_{1}$of the clusters (line 6 in Alg 1). Within the retained clusters, we perform weighted sampling from sub-clusters, prioritizing regions of the retained clusters which the model finds most challenging (line 7 in Alg. 1). Sub-clusters with higher $PPL$ are assigned greater sampling weights, enabling a hierarchical focus on the most critical regions. For instance, during Llama-2 training, GitHub code emerged as the most challenging cluster. Within this cluster, the Focus stage further emphasizes sampling from C++ code, which proved more difficult for the model, over Python code. In this phase of training, we restrict the weighted sampling of data points to this reduced subset for $p_{2}$ steps. $s_{1}$ and $p_{2}$ are user-controlled hyperparameters.
    2. (b)
      Multi-epoch training: We discard $s_{1}$of the data blocks (line 10 in Alg. 1) with the lowest PPL values. In doing so, we provide a mechanism for shifting the model’s focus towards learning data blocks that were determined to be difficult.
@@
    Review: Next, we reintroduce all of the removed data blocks and train the model by randomly sampling from the entire corpus for $p_{3}$ steps (line 13 in Alg. 1). This ensures that we allow the model to review and revisit data blocks which it may have forgotten.
Algorithm 1 LFR Training Methodology
1:Training dataset $D$, model $M$ with initial parameters $\theta_{0}$, hyperparameters $p_{1}$, $s_{1}$, $p_{2}$, $p_{3}$, $reps$, and $epochs$.
2:Minimization of Equation 3.
3:$PPLs,\theta_{p_{1}}\leftarrow\textbf{Learn}(\theta_{0},D,p_{1})$
4:for $r=1,2,\ldots,reps$ do
5:     if $epochs==1$ then
6:         $D_{k}\leftarrow Cluster(D)$
7:         $Sort(PPLs,D_{k})$
8:         $S_{sub}\leftarrow(1-s_{1})\times D_{k}$
9:         $S_{1}\leftarrow sample(S_{sub},PPLs)$
10:     else
11:         $Sort(PPLs,D)$
12:         $S_{1}\leftarrow(1-s_{1})\times D$
13:     end if
14:     $\theta_{p_{2}}\leftarrow\textbf{Focus}(\theta_{p_{1}},S_{1},p_{2})$
15:     $PPLs,\theta_{p_{3}}\leftarrow\textbf{Review}(\theta_{p_{2}},D,p_{3})$
16:end for
$Return\ \theta$
Our training strategy is simple, intuitive and human-like. It gives the model an opportunity to learn from all of the data, prioritize and relearn forgotten data points, and review data samples from harder regions of the dataset more frequently than they would have been using random sampling. While the static clustering-based techniques Tirumala et al. (2023); Abbas et al. (2023); Kaddour (2023) presented in Section 2 allow for accelerated training, they are not designed to suit the multi-descent training dynamics observed in Section 4 and require pretrained model embeddings to calculate distance metrics. Furthermore, prior methods including perplexity-based pruning methods Marion et al. (2023) are static. Sections A.4 and A.5 characterize the data blocks found easy and hard by the LLM, and demonstrate why static, clustering-based data selection methods achieve poor downstream task performance. Lastly, our approach does not require any pretrained models to obtain embeddings. Our focused training strategy allows the model to absorb harder information (data blocks with higher perplexity) faster, by presenting them more number of times.
5 Evaluation
In this section, we present a comprehensive evaluation of LFR. We pretrain the Llama-2 models of sizes 120M, 300M, and 500M, and the GPT-2 models Radford et al. (2019) of various sizes between 124M and 1.5B parameters. We use the SlimPajama Soboleva et al. (2023) and OpenWebText dataset ope and train from scratch using 4 NVIDIA A6000 GPUs (Llama-2), 4 AMD MI210 GPUs (GPT-2 345M-1.5B), 4 AMD MI100 GPUs (GPT-2 124M). Our pretraining experiments utilize a fully distributed data parallel (DDP) approach. All model configurations and training hyperparameters of our experiments are detailed in Section A.1.
We also evaluate our pretrained models on various downstream tasks from different domains in Sections 5.3 and 5.4. Section 5.5 demonstrates the impact of the Focus and Review stages. Sections A.2 presents examples of data points from the SlimPajama dataset which were identified as easy and challenging by LFR. Sections A.5 and  A.4 presents an analysis on the data blocks dropped and retained by LFR in different phases of training across the different GPT-2 models. Together Sections A.2,  A.5, and  A.4 demonstrate that LLMs learn instructions and code after facts and anecdotes as illustrated in Figure 1. Finally, we conduct a sensitivity study in Section A.3, where we test LFR with different hyperparameters.
Model Tokens Arc_C Arc_E Boolq HS OBQA Piqa WG
120M-RS 50B 23.29 39.06 53.15 34.30 30.50 62.58 51.54
Pythia-160M 300B 20.1 44.0 40.0 35.82 29.59 61.80 49.7
120M-LFR 9.8B 23.61 39.52 54.86 35.44 30.56 63.21 53.88
300M-RS 50B 25.7 41.9 54.0 36.4 33.3 64.8 52.4
Pythia-410M 300B 30.3 47.1 55.3 40.1 36.2 67.2 53.4
300M-LFR 9.8B 36.61 44.52 54.86 48.44 40.56 66.21 56.88
500M-RS 50B 27.5 43.7 53.7 36.5 32.6 65.1 52.2
Pythia-1.0B 300B 37.05 56.99 60.83 47.16 39.40 70.70 58.70
500M-LFR 9.8B 38.81 53.10 58.72 50.65 43.10 68.66 58.72
Table 2: Zero-shot performance (acc_norm for all except Winogrande and Boolq which use acc) on Common Sense Reasoning Tasks evaluated using LLM Evaluation Harness Gao et al. (2024). RS refers to the random sampling baseline, HS refers to HellaSwag, and WG refers to Winogrande. The model with the highest performance (measured by acc_norm) is highlighted in bold. Notably, LFR models are trained using only 3.2% of the tokens required to train Pythia models of comparable size, yet they achieve higher accuracy in 55.55% of cases. Additionally, LFR models consistently outperform the random sampling baseline by a large margin, despite being trained on 19.6% of the pretraining tokens.
5.1 LFR Configuration
We pretrain the Llama models for a total of 100k steps and train on 9.8B tokens based on the Chinchilla scaling law Hoffmann et al. (2022). All training hyperparameters are described in Section A.1. First, we Learn for 20k steps ($p_{1}=20k$). Next, we cluster the data and discard 57.2% of the clusters, retaining only the 3 most challenging clusters out of 7 based on their $PPL$ values ($s_{1}=50$). We then apply the Focus stage for 60k steps ($p_{2}=60k$), prioritizing the retained high-PPL clusters. The clustering results using ($k=7$) are shown in Figure 6. It takes <10min to cluster which can be hidden by the high training latency. We provide a detailed analysis on the hierarchical clustering and the data points found easy and difficult in Section 5. Lastly, we Review the entire dataset for the last 20k steps ($p_{3}=20k$). In the case of the GPT-2 models, we Learn for 1 epoch ($p_{1}=1$), Focus on 50% of the data for 1 epoch ($s_{1}=50,p_{2}=1$), Review the entire dataset for another epoch ($p_{3}=1$), and Focus on 30% of the data for 5 epochs ($reps=2,s_{2}=70,p_{4}=1$).
We chose this configuration based on our limited pretraining budget and profiling in Section 3.2, which showed that 78.5% and 25% of data samples are forgotten at least once during training for Llama and GPT-2 models, respectively. Figure 3 reveals that forgotten samples are typically forgotten multiple times, requiring an average of 4 presentations to be learned. For GPT-2, we use the first three phases to identify these samples and allocate 5 epochs focusing on 30% of them in the final phase to ensure long-term retention. Additionally, all samples are reviewed through Phases 1 and 3, and Section A.3 evaluates LFR’s sensitivity to hyperparameters $p_{1}$, $s_{1}$, $p_{2}$, $p_{3}$, and $reps$.
5.2 Baselines